Grade 12Wave Motion

The number of possible natural oscillation of air column in a pipe closed at one end of length 85 cm whose frequency lie below 1250hz are...{v=340m/s}

Aniket
8 Years agoGrade 12

2 Answers

Arun
8 Years ago
Dear Aniket
In fundamental mode
λ/4 = 0.85
λ = 4*0.85
f = v/λ
f = 340/4*0.85 = 100 Hz
Therefore, possible frequencies = 100 Hz, 300 Hz, 500 Hz, 700 Hz, 900 Hz,1100 Hz below 1250 Hz.
Hence total frequencies = 6
Regards
Arun (askIITians forum expert)
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

In fundamental mode
λ/4 = 0.85
λ = 4*0.85
f = v/λ
f = 340/4*0.85 = 100 Hz
Therefore, possible frequencies = 100 Hz, 300 Hz, 500 Hz, 700 Hz, 900 Hz,1100 Hz below 1250 Hz.
Hence total frequencies = 6

Thanks and Regards